Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Urumqi (yang wen jun at Sun Mar 6 08:07:05 2005) Success :
because the source is outside dish limits, so the experiment have to stop at UT 07:53,

Onsala (Michael Lindqvist at Sun Mar 6 13:37:02 2005) Success :
Source to low elevation last scans. Otherwise no known problems.

Hartebeesthoek (Jonathan Quick at Sun Mar 6 14:06:02 2005) Minor failures : RFI
Intermittent RFI throughout experiment. No other known problems.

Torun (K. Borkowski at Mon Mar 7 09:43:17 2005) Minor failures : LATE
Start late by about 15 min (the first 2 scans lost). No other known problems.

Jodrell_Bank (Alastair Gunn at Mon Mar 7 12:30:03 2005) Success :
MkII was used. Lovell telescope drive failed completely just prior to this experiment. Schedule was run using the Mk2 telescope instead. No known problems.

Cambridge (Alastair Gunn at Mon Mar 7 12:30:32 2005) Success :
No known problems.

Westerbork (Tony Foley at Mon Mar 7 15:23:41 2005) Minor failures : WIND HIGH
Wb used in phased array mode. Still snow in telescopes We used a tied array with telesopes 012356789 (RT4 out with oscillation and pointing problems) From 08:15-09:11 RT0 out with pointing problems.

Effelsberg (A. Kraus at Fri Mar 11 12:43:10 2005) Minor failures : WIND
Stopped due to snowfall: UT 1030-1330.